OpenStudy (anonymous):

how can (sqrt (88))/2 be the same as sqrt(22)?

OpenStudy (anonymous):

88 is 4 times 22. That's how

OpenStudy (agent47):

sqrt(88)/2 = sqrt(4*22)/2 = 2*sqrt(22)/2 = sqrt(22)

OpenStudy (anonymous):

thanks now I get it